OpenStudy (this):

can some one plllz help me -9p^2+14-1=8p A,0;one real solution B,0; two imaginary solutions C,72;two real solution D,0;two real solution

hero (hero):

Are you sure you posted the problem correctly? Are you sure it isn't -9p^2 + 14p - 1 = 8p

OpenStudy (this):

thos is right -9p^2 + 14p - 1 = 8p

hero (hero):

That's what I figured.

hero (hero):

Okay, so subtract 8p from both sides to get -9p^2 + 6p - 1 = 0

hero (hero):

We can use b^2 - 4ac for guidance

OpenStudy (this):

yes

hero (hero):

6^2 - 4(-9)(-1) 6^2 + 36(-1) 6^2 - 36 36 - 36 0 So at least we know we won't have to deal with any imaginary solutions. And since we end up with zero, that pretty much means only one solution.

OpenStudy (anonymous):

-9p^2 + 14p - 1 = 8p 9p^2 - 14p +8p + 1 = 0 9p^2 -6p + 1 = 0

OpenStudy (anonymous):

b^2 - 4ac 36 - 36 = 0 So...it has Two equal roots

hero (hero):

If they are equal, then that means one solution bro.

OpenStudy (anonymous):

So...u can Take it has one...

OpenStudy (anonymous):

Yup..)...i was coming there @Hero

OpenStudy (this):

thanks @Hero